OpenStudy (stuck_in_the_middle):

a) an inductor of 85 mH is connected in series with 100 Ω resistor. If a sinusoidal current of 25 mA at 50 Hz flows in the circuit, determine: a) the voltage dropped across the inductor; b) the voltage dropped across the resistor; c) the impedance of the circuit; d) the supply Voltage; e) the phase angle. b) A capacitor of 22 µF is connected in series with a 640 Ω resistor if a sinusoidal current of 12 mA at 50 Hz flows in the circuit, determine: a) the voltage dropped across the capacitor; b) the voltage dropped across the resistor; c) the impedance of the circuit; d) the supply v

OpenStudy (stuck_in_the_middle):

e) the phase angle. (c) A series circuit comprises of an inductor of 85 mH, a resistor of 220 Ω and a capacitor of 22 µF. If a sinusoidal current of 40 mA at 50 Hz flows in this circuit, determine: a) the voltage developed across the inductor; b) the voltage dropped across the capacitor c) the voltage dropped across the resistor; d) the impedance of the circuit; the supply voltage; e) the phase angle.

OpenStudy (stuck_in_the_middle):

Any help would be deeply appreciated.

OpenStudy (ybarrap):

a) $$ \large{ i(t)=.025\sin 2\pi ft\\ =.025\sin 2\pi50 t\\ =.025\sin 100\pi t\\ v_L(t)=L\cfrac{d}{dt}i(t)=L(0.025\times100\pi)\cos (100\pi t)\\ =.085\times2.5\pi\cos (100\pi t)\\ \approx 0.67\cos (100\pi t)~Volts\\ } $$ b) $$ \large{ v_R(t)=i(t)R\\ =100\times.025\sin 2\pi ft\\ =2.5\times \sin 100\pi t~Volts\\ } $$ c) $$ \large{ Z(t)=\cfrac{v(t)}{i(t)}\\ =\cfrac{v_L(t)+v_R(t)}{.025\sin 100\pi t} } $$ Substitute values determined in parts a and b above d) $$ \large v(t)=v_L(t)+v_R(t) $$ e) Use trig identities to simplify \(v(t)\) and put into form: \(A\cos (\omega t+\phi)\). Where \(\phi\) is the phase angle. Next circuit is handled similarly but will require integration rather than differentiation.

OpenStudy (stuck_in_the_middle):

Thank you ybarrap

OpenStudy (ybarrap):

you're welcome